OpenStudy (andrewkaiser333):

which of the following could be the equation of a line parallel to he line y=4x-7?

OpenStudy (anonymous):

y=4x+(any other number besides 7)

OpenStudy (andrewkaiser333):

y=(1/4)x-7 y=(-1/4)x-7 y=-4x+3 or y=4x+3

OpenStudy (***[isuru]***):

when a line is parallel to another line the slope of both lines r equal... and when the lines r given in general format y = mx + c m or the coefficient of x is the slope... so in ur question y= 4x -7 4 is the slope... so u have to chose any other equation with the slope 4

OpenStudy (***[isuru]***):

in other words any equation with 4x included :)

OpenStudy (andrewkaiser333):

is is the last answer i put correct ????

OpenStudy (anonymous):

ya it is

OpenStudy (andrewkaiser333):

kk

OpenStudy (andrewkaiser333):

can you give isuru a medal for mee i gave you one

OpenStudy (anonymous):

Only your last answer was though

OpenStudy (andrewkaiser333):

kk